OpenStudy (anonymous):

7 sides A) two thousand eight hundred eighty degrees 12 sides B) nine hundred degrees 14 sides C) one thousand eight hundred degrees 18 sides D) two thousand one hundred sixty degrees match please

OpenStudy (texaschic101):

sum of interior angles of polygon : 180(n - 2).....n being the number of sides 7 sides.....180(n - 2) = 180(7 - 2) = 180(5) = 900 degrees ... (B) 12 sides...180(n - 2) = 180 (12 - 2) = 180(10) = 1800 degrees (C) 14 sides...180(n - 2) = 180(14 - 2) = 180(12) = 2160 degrees (D) 18 sides...180(n - 2) = 180(18 - 2) = 180(16) = 2880 degrees (A) any questions ?